force Fx acts on a particle that has a mass of 2 kg. The force is related to the position x of the particle by the formula F 3, where C-0.5 if x is in meters and Fy is in Newton (a)What are the SI units of C? (Use the following as necessary: N and m.) (b) Find the work Wdone by this force as the partidle moves from x - 2.8 m to x - 1.6 m (C) At x= 1.6 m the force points opposite the direction of the particles velocity (speed is 12 m/s), what is its speed at 2.8 m? m/s Can you tell its direction of motion at x-2.8 m using only the work-kinetic energy theorem yes O no Explain. has not been graded yet
Answer #1

a)

Fx = C x3

N = C [m3]

C = Nm-3

so unit are Newton/meter3

b)

work done is given as

W = \int_{x_{i}}^{x_{f}} F dx

given that xi = 2.8 , xf = 1.6 m

W = \int_{2.8}^{1.6} C x3 dx

W = (C/4) (1.64 - 2.84)

W = (0.5/4) (1.64 - 2.84)

W = - 6.86 J

c)

vi = initial speed at 1.6 m, = 12 m/s

vf = final speed = ?

using work-change in kinetic energy

W = (0.5) m (vf2 - vi2)

6.86 = (0.5) (2) (vf2 - (12)2)

vf = 12.3 m/s